OpenStudy (anonymous):

Solve for x: |x| − 8 = −5 x = −13 and x = −3 x = 3 and x = −3 x = 3 and x = 13 No solution i thonk its B.. am i right?

Mehek (mehek14):

let's see 3 - 8 = -5 -3 - 8 = -11

OpenStudy (anonymous):

so its no solution ? @Mehek14

Mehek (mehek14):

the only solution there is is 3

OpenStudy (anonymous):

oh ok... thankyou!

Directrix (directrix):

@Madisons_xo_beauty On first pass through the problem, I got 3 and -3. Do you want to go through this again or are you good with the answer of whatever you already have?

OpenStudy (anonymous):

ya i think im good, but thanks @Directrix

Directrix (directrix):

What is your answer, if I may ask. I question this >>> the only solution there is is 3

Directrix (directrix):

|x| − 8 = −5 |x| = -5 + 8 |x| = 3 x = ? or ? @Madisons_xo_beauty